The pressure below the depth of 25 m of a homogenous fluid is 50.473 psi.


Calculate specific weight in kN/m3,density of fluid in kgm/m3 and its specific


gravity

1
Expert's answer
2022-02-11T08:47:00-0500

Given:

"h=25\:\rm m"

"p=50.473\: \rm psi=347999\: Pa"


The specific weight

"g\rho=\frac{P}{h}\\
=\frac{347999}{25}=13920\:\rm N/m^3=13.92\: kN/m^3"

The density of fluid

"\rho=\frac{P}{gh}=\frac{13920}{9.8}=1420\:\rm kg/m^3"

The specific gravity

"\rho/\rho_w=1420/1000=1.42"